# LoRAMoE: Revolutionizing Mixture of Experts for Maintaining World Knowledge in Language Model Alignment

### 18 Dec 2023 v2

@@ 3.2.3 Localized Balancing Constraint
Formally, we define the importance matrix $\mathbf{Q}$ of LoRAMoE layer and $\mathbf{Q}_{n,m}$ denotes the sum of router values of the $n$-th expert for the $m$-th training sample in a batch, which can be represented as follows:
@@
where $N$ and $T_{m}$ denote the number of experts and the number of tokens of $m$-th training sample, respectively. $x_{j}$ is the hidden input of the $j$-th token to the LoRAMoE layer. We then define the coefficient matrix $\mathbf{I}$ with the same size of $\mathbf{Q}$, corresponding to the importance matrix $\mathbf{Q}$. $\mathbf{I}_{n,m}$ denotes the importance coefficient of $\mathbf{Q}_{n,m}$, which can be written as follows:
$\mathbf{I}_{n,m}=\left\{\begin{array}[]{lr}1+\delta,&\text{Type}_{e}(n)=\text{Type}_{s}(m).\\ 1-\delta,&\text{Type}_{e}(n)\neq\text{Type}_{s}(m).1\end{array}\right.$ (14)
where $\delta\in[0,1]$ controls the degree of imbalance between experts groups. $\text{Type}_{e}(n)$ and $\text{Type}_{s}(m)$ are pre-defined target type of $n$-th expert and the task type of $m$-th training sample in a batch, respectively. We categorize the instruction data into two distinct groups: CBQA and other downstream task data. The CBQA serves to enable expert groups to align human instructions with knowledge, whereas the remaining data is allocated to task-focused experts to boost task performance.
Specifically, suppose that $\mathbf{I}_{i,k}$ and $\mathbf{I}_{j,k}$ denote the importance coefficient of the $i$-th and $j$-th expert for the $k$-th sample, respectively. Two experts intentionally assigned to the same panel hold equivalent values at corresponding positions in the coefficient matrix (i.e., $\mathbf{I}_{i,k}=\mathbf{I}_{j,k}$). This implies that their importance is assigned equal weight. On the contrary, two experts in distinct groups possess divergent values at corresponding positions in their coefficient matrix (i.e., $\mathbf{I}_{i,k}\neq\mathbf{I}_{j,k}$).
We define the localized balancing constraint loss $\mathcal{L}_{lbc}$ to quantify the dispersion of the weighted importance matrix $\mathbf{Z}=\mathbf{I}\circ\mathbf{Q}$, which can be mathematically represented as:
@@
where $\sigma^{2}(\mathbf{Z})$ and $\mu(\mathbf{Z})$ represent the variance and mean of $\mathbf{Z}$, respectively. Specifically, for a specific sample, given that experts on the same panel possess identical values in the coefficient matrix $\mathbf{I}$, the reduction of the optimized loss $\mathcal{L}_{lbc}$ results in a progressive equalization of their importance. On the other hand, experts engaged in distinct panels are assigned differing values in $\mathbf{I}$, thus, the convergent state of the loss essentially equates to attaining a soft-weighted balance among these different groups with the $\mathbf{I}$ scaling the $\mathbf{Q}$. In this state, the router intensifies its focus on the particular expert group according to the type of task, yet refrains from fully obscuring other groups, thereby preserving the capacity for generalization.
Ultimately, optimizing $\mathcal{L}_{lbc}$ reduction achieves a balance in the importance of experts within the same group, while also ensuring different groups concentrate on different capabilities. The overall loss of LoRAMoE can be represented as follows:
@@
where $\mathcal{L}$ is the next-token prediction loss of large language models and $\mathcal{L}_{lbc}$ is the localized balancing constraint loss for all LoRAMoE layers. $\beta$ controls the strength of localized balancing constraint.
During the fine-tuning phase, we freeze the base model and the trainable parameters are those of the experts and routers within the LoRAMoE layers. LoRAMoE significantly conserves resources compared to fine-tuning the full range of parameters of models. In the inference process, the router autonomously determines and assigns output weights to all experts, thereby obviating the need for pre-specified data types.
4 Experiments
4.1 Experiment Setup
In this section, we introduce the training implementation for LoRAMoE. We only replace the linear layer in the feed-forward neural network of LLM with the LoRAMoE layer. We initialize each LoRAMoE layer with six experts, of which three experts are dedicated to addressing downstream tasks, and the other three are responsible for aligning world knowledge in the base model with the human instructions. The hyperparameters for control constraint strength $\beta$ and degree of imbalance $\delta$ are both set to $0.1$. For the experts based on low-rank adapter, the $\alpha$, and $r$ are set to $32$ and four, respectively. The dropout is $0.05$, and the learning rate for both the experts and router in the LoRAMoE layer is $42e-$ 4. The training dataset is the 3 million set the same as the one described in Section 2.1. We freeze the parameters of the base model, rendering only the experts and router in LoRAMoE trainable. The global batch size is set to $64$, and all our experiments were conducted on 32-card A100 80G.
4.2 Results
Task baseline
@@ 4.2 Results
Table 2 displays the performance of LoRAMoE with 3 million training samples and compares this result with the outcomes of directly applying SFT to the model or utilizing LoRA tuning. We report the results on the same test set as discussed in Section 2.1. The results show that the language model with LoRAMoE gets good performance on both world knowledge benchmarks and others, indicating the effectiveness of avoiding knowledge forgetting while improving various tasks.
For world knowledge benchmarks, first of all, the catastrophic collapse observed in Section 2 did not occur. On the contrary, the model with the LoRAMoE plugin even outperforms the one fine-tuned solely with the CBQA dataset, such as on TriviaQA and HotpotQA. Besides, compared to vanilla SFT using the same amount of data, LoRAMoE achieves up to a 64% improvement in world knowledge benchmarks, with an average increase of 35%. By comparing the performance of LoRAMoE with and without the implementation of the local balance expert loss (the last two columns in Table 2), we observed that $\mathcal{L}_{lbc}$ can facilitate the performance on the world knowledge benchmarks.
LoRA tuning is intuitively a method involving smaller degrees of parameter modification than vanilla SFT (He et al., 2021a). However, it is observed that the forgetting of world knowledge still occurs. Its performance in CBQA is generally lower than the baseline, and there is an average decrease of 22% compared to direct single-task tuning. Multiple parallel experts are helpful in achieving excellent results in balancing world knowledge retention and the scaling up of fine-tuning data.
@@
For other downstream tasks, LoRAMoE is capable of achieving performance close to or even surpassing that of direct SFT, For instance, in all reading comprehension tasks (i.e., Race, ReCoRD, multiRC), LoRAMoE achieved superior performance. Besides, $\mathcal{L}_{lbc}$ improves outcomes for LoRAMoE in the vast majority of tasks, both world knowledge benchmarks and others. Notably, for reading comprehension, NLI, and the original CBQA dataset, the benefits of this method were quite substantial, up to 11.8%. This indicates capability partitioning in the expert group benefits the performance. Further, LoRAMoE holds significant promise for multi-task learning.
@@ 4.3 Visualizing the experts utilization
Figure 7: Visualization of routers’ weight on different types of data, where Group 1 refers to the experts dedicated to aligning the world knowledge in the base model with the human instruction and Group 2 refers to the experts focus on downstream tasks. The utilization rate of the group of experts diverged significantly across tasks.
To confirm the effectiveness of LoRAMoE in specializing the grouped experts, we visualize their weight assigned by the router when encountered with data from downstream tasks and knowledge benchmarks respectively, as illustrated in Figure 7.
There is a distinct contrast in the utilization of the two expert groups when dealing with world knowledge benchmarks and other downstream tasks. This suggests that the routers can automatically allocate specific tasks to experts with corresponding abilities during the inference phase. Specifically, the expert group requested to prioritize aligning parametric knowledge with human instructions is greatly employed in world knowledge benchmarks (e.g., TriviaQA, Natural Question, and HotpotQA), underscoring their vital role in preventing world knowledge forgetting. This corresponds to the fact we state in Section 2 that supervised fine-tuning boosts the model’s capabilities in these tasks by associating pre-stored world knowledge in the model with human instructions. On the other hand, experts assigned to focus on enhancing performance in downstream tasks are given increased prominence when encountering these tasks. Through this visualized result, we find that some downstream tasks still require another group of experts. It is reasonable. For example, in reading comprehension tasks, the knowledge learned by the model during pre-training can better assist in making factual judgments. This phenomenon is even more pronounced in language-based tasks. In the WSC task (Levesque et al., 2012), the router allocates an average of about 45% of its attention to the expert group responsible for world knowledge.
@@ 5 Related Work
Parameter-Efficient Fine-tuning. With the significant rise in the number of parameters in language models, parameter-efficient fine-tuning (PEFT) (He et al., 2021a) has become an important research trend. It can consume less resources while fine-tuning the large language models. Several methods achieve more efficient fine-tuning of the language model, including prompt LoRA (Hu et al., 2021), adapters (Houlsby et al., 2019), and prompt learning (Lester et al., 2021). PEFT based on Lora-rank adapters (Hu et al., 2021) is widely used, which introduce two trainable low-rank matrices for each fully connected layer, to achieve significant savings in training resources without adding additional inference computation cost. In this paper, We introduce low-rank to the expert networks of LoRAMoE to reduce fine-tuning consumption, which increases the number of experts to improve performance while not significantly increasing computational resources.
Mixture-of-Experts. Mixture-of-Experts (MoE) (Jacobs et al., 1991) modifies the feed-forward neural network layer to sparsely activated experts, which significantly enlarges the model without remarkably increasing the computational cost. The exploration of MoE has attracted more and more attention in recent years, including the early sample-level MoE (Jacobs et al., 1991) up to the token-level MoE (Shazeer et al., 2016; Lepikhin et al., 2020; Du et al., 2022; Riquelme et al., 2021; Fedus et al., 2021; Xue et al., 2022) that have become mainstream nowadays. Meanwhile some researchers (Zhou et al., 2022; Chi et al., 2022) aim to investigate the router selection problem in MoE. However, the vast majority of the works are trying to significantly increase the model parameters and reduce the computational cost. Differently, Chen et al. (2023) explores the advantages of MoE-based language modeling in continuous learning. Our approach is significantly different. We use a MoE-like structure to address the parametric knowledge retention issue in LLMs, rather than significantly expanding the count of model parameters.
Multi-LoRA Architecture. Some researchers have also proposed the utilization of multiple LoRAs to improve the performance of the models or to be more advantageous in some aspects. Huang et al. (2023) proposed LoraHub which trains several LoRAs and chooses different LoRA combinations based on data types during the inferencing phase. MOELoRA (Liu et al., 2023) fine-tuned language models through the incorporation of MoE structures, thereby boosting their efficacy in multitasking within the medical domain. However, these methods take the data type as the input of the router during training, which necessitates prior knowledge of the data type during inference, i.e., choosing the combination of LoRAs based on the data type. Zadouri et al. (2023) introduced Mixture of Vectors and Mixture of LoRA to reduce the resource consumption of fine-tuning the large language models and improve the performance on unseen tasks. Sheng et al. (2023) proposed S-LoRA, a system that can serve thousands of LoRA adapters from a single machine. These methods are very different from LoRAMoE. We explored for the first time that the expansion of instruction data seriously damages the knowledge of the pre-trained language models during the fine-tuning phase. We further propose to address conflicts through the MoE structure and to utilize the LoRA to reduce resource consumption. Simultaneously LoRAMoE is also an end-to-end approach that does not require a priori knowledge of data types for inference.
@@ 6 Conclusion
In this paper, we delve into the consequences of increasing data quantity on the instruction task during the supervised fine-tuning stage. We uncover that an extensive expansion in data quantity can critically impair the world knowledge of LLMs, consequently causing world knowledge forgetting. To address this conflict, we then introduce LoRAMoE, a new MoE-based trainable LLM plugin that separates networks concentrate on different capabilities with localized balancing constraints. Extensive experiments have shown that LoRAMoE is adept at resolving this conflict. It preserves the world knowledge inside the LLMs, while simultaneously enhancing the performance of other downstream tasks within the instruction data during the fine-tuning process.
Appendix A Appendix
A.1 Details about fine-tuning datasets
Table 3 shows the specific tasks covered by the 3-million-sample dataset we used in Section 2 and the amount of data for each task. The five million fine-tuning data we use includes three million versions and their variants from data augmentation strategies. The 1-million-sample version is the subset of the original 3-million-sample dataset.
@@ A.1 Details about fine-tuning datasets
TriviaQA (Han et al., 2019)
78785 254
closed-book QA
Natural Questions (Kwiatkowski et al., 2019)
104071 357
closed-book QA
Hotpot QA (Qi et al., 2019)
72798 5622
closed-book QA
WSC (Levesque et al., 2012)
554 146
coreference resolution
WinoGrande (Sakaguchi et al., 2021)
40398 1767
coreference resolution
Flores (Guzmán et al., 2019)
0 1600
machine translation
WMT  [^2] https://huggingface.co/datasets/wmt14, https://huggingface.co/datasets/wmt16
500000 -
machine translation
RTE  [^3] https://aclweb.org/aclwiki/Recognizing_Textual_Entailment
2490 3000
NLI
ReCoRD (Zhang et al., 2018)
100730 10000
reading comprehension
AX-g  [^4] https://super.gluebenchmark.com
0 356
NLI
multiRC (Khashabi et al., 2018)
27243 9693
reading comprehension
anli r1/r2/r3 (Liu et al., 2020)
162874 -
NLI
qqp (Wang et al., 2017)
363846 -
NLI
Xsum (Narayan et al., 2018)
204045 11334
single-document summarization
Race (Lai et al., 2017)
87866 4934
reading comprehension
duorc-selfRC (Saha et al., 2018)
60721 -
reading comprehension
AG-news (Zhang et al., 2015)
120000 -
topic classification
yelp review (Zhang et al., 2015)
650000 -
sentiment classification
@@
openai tldr  [^5] https://github.com/openai/summarize-from-feedback
232188 -
summarization
@@
Table 3: Details about the tasks in our fine-tuning dataset. ”-” means we do not use the test set of this dataset for evaluation.

### 05 Mar 2024 v3

@@ 3.2 Localized Balancing Constraint
Formally, we define the importance matrix $\mathbf{Q}$ of the LoRAMoE layer and $\mathbf{Q}_{n,m}$ denotes the sum of router values of the $n$-th expert for the $m$-th training sample in a batch, which can be represented as follows:
@@
where $N$ and $T_{m}$ denote the number of experts and the number of tokens of $m$-th training sample, respectively. $x_{j}$ is the hidden input of the $j$-th token. We then define the coefficient matrix $\mathbf{I}$ with the same size of $\mathbf{Q}$, corresponding to the importance matrix $\mathbf{Q}$. $\mathbf{I}_{n,m}$ denotes the importance coefficient of $\mathbf{Q}_{n,m}$, which can be written as follows:
$\mathbf{I}_{n,m}=\left\{\begin{array}[]{lr}1+\delta,&\text{Type}_{e}(n)=\text{Type}_{s}(m)\\ 1-\delta,&\text{Type}_{e}(n)\neq\text{Type}_{s}(m)\end{array}\right.$ (7)
where $\delta\in[0,1]$ controls the degree of imbalance between experts types. $\text{Type}_{e}(n)$ and $\text{Type}_{s}(m)$ are pre-defined target type of $n$-th expert and the task type of $m$-th training sample in a batch, respectively.
We categorize the instruction data into two distinct types: world knowledge-related tasks such as TriviaQA, and other downstream tasks such as Flores. Then, we enable a portion of experts to learn on world knowledge-related tasks to align human instructions with world knowledge, while making other experts focus more on enhancing the performance of downstream tasks. Formally, suppose that $\mathbf{I}_{i,k}$ and $\mathbf{I}_{j,k}$ denote the importance coefficient of the $i$-th and $j$-th expert for the $k$-th sample, respectively. If experts are in the same group, their values at corresponding positions in the coefficient matrix are identical, i.e., $\mathbf{I}_{i,k}=\mathbf{I}_{j,k}$. This indicates that these experts have the same importance because they are assigned to focus on learning the same type of tasks. On the contrary, the values of experts from distinct groups at their coefficient matrix are different, i.e., $\mathbf{I}_{i,k}\neq\mathbf{I}_{j,k}$.
The localized balancing constraint loss $\mathcal{L}_{lbc}$ is defined to measure the dispersion of the weighted importance matrix $\mathbf{Z}=\mathbf{I}\circ\mathbf{Q}$, which can be mathematically represented as:
@@
where $\sigma^{2}(\mathbf{Z})$ and $\mu(\mathbf{Z})$ represent the variance and mean of $\mathbf{Z}$, respectively. Specifically, if a specific sample is from the world knowledge-related dataset, experts focusing on solving this type will have larger values in the coefficient matrix $\mathbf{I}$. Optimizing the loss $\mathcal{L}_{lbc}$ reducing can make corresponding experts learn more from this sample and be assigned a larger weight by the router. Meanwhile, experts solving the same type of task are balanced such as Shazeer et al. (2016). In addition, the constraint is soft to encourage cooperation among experts to preserve the capacity for generalization.
Overall, localized balancing constraint $\mathcal{L}_{lbc}$ achieves a localized balance between two types of experts: one specializes in leveraging world knowledge by training more on world knowledge-related datasets, while the other concentrates on various downstream tasks. The loss of LoRAMoE can be represented as follows:
@@
where $\mathcal{L}$ is the next-token prediction loss of LLMs and $\beta$ controls the strength of localized balancing constraint. In the training phase, we freeze the backbone model and the trainable parameters are only those of the experts and routers within the LoRAMoE layers. In the inference process, the router automatically assigns weights to all experts, which avoids the need for pre-specified data types.
Task Name Baseline
@@
4 Experiments
4.1 Experiment Setup
In this section, we introduce the training implementation for LoRAMoE. We only replace the linear layer in the feed-forward neural network of LLM with the LoRAMoE layer, initializing each layer with six experts, of which three experts are dedicated to addressing downstream tasks, and the other three are responsible for leveraging world knowledge in the base model by learning on its related tasks. The hyperparameters for control constraint strength $\beta$ and degree of imbalance $\delta$ are both set to $0.1$. For LoRA settings, the $\alpha$, and $r$ are set to $32$ and four for the main result, respectively. The dropout is $0.05$, and the learning rate is $42e-$ 4. The training dataset is the 3 million set the same as the one described in Appendix A, so as the evaluation settings. We freeze the parameters of the base model, rendering only the experts and router in LoRAMoE trainable. The batch size per node is set to $16$.
4.2 Main Results
@@
Table 2 displays the performance of LoRAMoE and compares this result with the outcomes of directly applying SFT to the model or utilizing LoRA tuning. The results show that the language model with LoRAMoE gets good performance on both world knowledge benchmarks and others, indicating its effectiveness in avoiding knowledge forgetting while improving multi-tasking abilities.
For world knowledge benchmarks, contrary to the catastrophic collapse seen in Section 2, LoRAMoE not only avoids this issue but also surpasses the model fine-tuned solely with the CBQA dataset. LoRAMoE shows a significant performance boost on world knowledge benchmarks over vanilla SFT, with up to a 63.9% improvement and an average increase of 35.3%.
For other downstream tasks, LoRAMoE is capable of achieving performance close to or even surpassing that of direct SFT. For instance, in all reading comprehension tasks (i.e., Race, ReCoRD, multiRC), LoRAMoE achieved superior performance.
We also compare our method against PEFT by single LoRA. The knowledge forgetting also occurred during the single LoRA-tuning, as it is essentially the same as vanilla SFT (Hu et al., 2021). Compared with a single LoRA, multiple collaborative LoRAs in LoRAMoE enhance both world knowledge retention and multitasking performance. They offer an average boost of 30.9% in world knowledge benchmarks and 8.4% in other downstream tasks.
@@
Besides, $\mathcal{L}_{lbc}$ improves outcomes for LoRAMoE in the vast majority of tasks, both world knowledge benchmarks and others. Notably, for reading comprehension, NLI, and the original CBQA dataset, the benefits of this method were quite substantial, up to 17.6%. This indicates capability partitioning in the expert group benefits the performance in multi-task learning.
# Experts
# LoRA
Rank
# Trainable
Param.
Avg.
Results
6 4 0.57% 58.21
4 4 0.38% 55.84
8 4 0.76% 56.58
6 8 1.07% 58.11
6 16 2.08% 58.86
Table 3: Performance of LoRAMoE varies with the number of experts and LoRA rank across all test sets. This includes the average results on both the world knowledge benchmark and all other downstream tasks. LoRAMoE shows stability to parameter changes.
4.3 Sensitivity Analysis
In this section, we analyze the parameter sensitivity of LoRAMoE. Keeping other settings constant, we vary the number of experts and the rank of LoRA. The average performance with varied parameter settings on all test sets including the world knowledge benchmark and all other downstream tasks is shown in Table 3. In Appendix D there are detailed results.
As the number of trainable parameters increases, performance is generally stable. the number of 6 experts is the most beneficial choice, as more experts do not lead to higher performance. While the increase in LoRA rank improves the model’s capabilities somewhat, it brings about an exponential rise in trainable parameters.
@@ 4.4 Visualizing the Experts Utilization
To confirm the effectiveness of LoRAMoE in specializing the experts with two types, we visualize their weight assigned by the router when encountered with data from downstream tasks and knowledge benchmarks respectively, as illustrated in Figure 7.
Figure 7: Visualization of routers’ weight on different types of data, where type 1 refers to the experts dedicated to aligning the world knowledge in the base model with the human instruction and type 2 refers to the experts that focus on downstream tasks. The utilization rate of the type of experts diverged significantly across tasks.
There is a distinct contrast in the utilization of the two types of experts when dealing with world knowledge benchmarks and other downstream tasks. This suggests that the routers can automatically allocate specific tasks to experts with corresponding abilities during the inference phase. Specifically, the experts requested to leverage world knowledge are greatly employed in world knowledge benchmarks (e.g., TriviaQA, Natural Questions, and HotpotQA), underscoring their vital role in preventing world knowledge forgetting. This corresponds to the fact we state in Section 2 that supervised fine-tuning boosts the model’s capabilities in these tasks by associating pre-stored world knowledge in the model with human instructions. On the other hand, experts assigned to focus on enhancing performance in downstream tasks are given increased prominence when encountering these tasks. Through this visualized result, we find that some downstream tasks still require experts of another type. It is reasonable. For example, in reading comprehension tasks, the knowledge learned by the model during pre-training can better assist in making factual judgments. This phenomenon is even more pronounced in language-based tasks. In the WSC task (Levesque et al., 2012), the router allocates an average of about 45% of its attention to the experts responsible for world knowledge.
@@ 5 Related Work
Parameter-Efficient Fine-tuning. With the size of language models growing larger, parameter-efficient fine-tuning (PEFT (He et al., 2021)) has become crucial for resource savings. Researchers have proposed several approaches such as LoRA (Hu et al., 2021), adapters (Houlsby et al., 2019), and prompt learning (Lester et al., 2021), to enhance fine-tuning efficiency. PEFT based on low-rank adapters (Hu et al., 2021) is popular and widely used, which introduces two trainable low-rank matrices in each fully connected layer, to achieve significant savings in training resources without adding additional inference computation cost. We apply low-rank techniques to the structure of experts to save resource consumption.
Mixture-of-Experts. The mixture of Experts (MoE) replaces the feed-forward neural network layer with sparsely activated experts, which significantly enlarges the model without remarkably increasing the computational cost (Jacobs et al., 1991). Currently, the token-level MoE architectures are widely used in pre-trained language models and vision models (Shazeer et al., 2016; Lepikhin et al., 2020; Du et al., 2022; Riquelme et al., 2021). In addition, researchers (Zhou et al., 2022; Chi et al., 2022) aim to investigate the router selection problem in MoE. Unlike these efforts to expand the model size and address the selection problem, we propose an MoE-style framework for multi-task learning and maintaining the world knowledge stored in LLMs.
Multi-LoRA Architecture. Researchers also have utilized multiple LoRAs for enhanced model performance. Huang et al. (2023) propose LoraHub to choose different LoRA combinations for task generalization. MOELoRA (Liu et al., 2023) leverage LoRA and MoE for task-specific tuning and multitasking, especially in healthcare. However, these methods need the data type as the input during the inference phase, which limits the application of the model to other tasks. Chen et al. (2023a) first introduces multiple LoRA serving systems and Sheng et al. (2023) proposes S-LoRA, a system that can serve thousands of LoRA adapters from a single machine. Chen et al. (2023b) introduces several experts to enhance the model’s ability for multimodal learning. Unlike these approaches, LoRAMoE introduces an MoE-style plugin and Localize Balancing Constraint to tackle world knowledge forgetting in LLMs, while enhancing the model’s ability to multi-task learning.
@@ 6 Conclusion
In this paper, we first delve into the conflict between improving LLM’s performance on downstream tasks by scaling up data during the SFT phase and discouraging world knowledge forgetting. To address this conflict, we then introduce LoRAMoE, a novelty framework for SFT, which introduces LoRAs as experts and integrates them by the router. Extensive experimental results demonstrate that LoRAMoE can foster collaboration among experts to enhance the model’s performance of downstream tasks, while preserving the world knowledge inside it.
7 Limitations
In this section, we discuss the potential limitations of our proposed method LoRAMoE. Firstly, although we have demonstrated the effectiveness of LoRAMoE in alleviating world knowledge forgetting while enhancing the downstream ability of the LLMs with SFT, we limit the model size to 7B due to resource and time constraints. Further work will be conducted on the larger LLMs, to understand the influence of large-scale SFT on these LLMs and to boost their multitasking abilities. Secondly, the localized balancing constraint can softly constrain the type of experts and balance the experts utilization. However, we haven’t studied the case where there are more experts types for a more fine-grained task category. Future work will be conducted on a more fine-grained understanding of the influence of SFT and the utilization of LoRAMoE.
Appendix A Details about Experiment Implementation
@@
TriviaQA (Han et al., 2019) 78785 254 closed-book QA
NQ (Kwiatkowski et al., 2019) 104071 357 closed-book QA
HotpotQA (Qi et al., 2019) 72798 5622 closed-book QA
WSC (Levesque et al., 2012) 554 146 coreference resolution
WinoGrande (Sakaguchi et al., 2021) 40398 1767 coreference resolution
Flores (Guzmán et al., 2019) 0 1600 machine translation
WMT  [^2] https://huggingface.co/datasets/wmt14, https://huggingface.co/datasets/wmt16 500000 - machine translation
RTE  [^3] https://aclweb.org/aclwiki/Recognizing_Textual_Entailment 2490 3000 NLI
ReCoRD (Zhang et al., 2018) 100730 10000 reading comprehension
AX-g  [^4] https://super.gluebenchmark.com 0 356 NLI
multiRC (Khashabi et al., 2018) 27243 9693 reading comprehension
anli r1/r2/r3 (Liu et al., 2020) 162874 - NLI
qqp (Wang et al., 2017) 363846 - NLI
Xsum (Narayan et al., 2018) 204045 11334 single-document summarization
Race (Lai et al., 2017) 87866 4934 reading comprehension
duorc-selfRC (Saha et al., 2018) 60721 - reading comprehension
AG-news (Zhang et al., 2015) 120000 - topic classification
yelp review (Zhang et al., 2015) 650000 - sentiment classification
@@
openai tldr  [^5] https://github.com/openai/summarize-from-feedback 232188 - summarization
@@
Table 4: Details about the tasks in our fine-tuning dataset. "-" means we do not use the test set of this dataset for evaluation.
Datasets. The seven tasks are closed-book question answering (CBQA), coreference resolution, natural language inference (NLI), abstract summarization, multi-lingual translation, reading comprehension, and text classification. Table 4 shows the composition of the 3-million-sample dataset. The five million fine-tuning data we use includes three million versions and their variants from data augmentation strategies. The 1-million-sample version is the subset of the original 3-million-sample dataset.
Evaluation. We utilize the opencompass [^6] https://opencompass.org.cn/ framework to run the evaluation process on the aforementioned tasks. Notably, considering previous work that has noted train-test overlap in CBQA datasets (Lewis et al., 2020), we elaborately select parts of the CBQA dataset without train-test overlap for our testing set, namely Filtered NQ and Filtered TriviaQA, to analyze the world knowledge of models better.
Appendix B The World Knowledge of LLM Further Declines after Being Trained with More Data
With the task types increasing, there is an inevitable trend to increase the amount of SFT training data. To further verify that a large-scale SFT training process can lead to knowledge forgetting of LLM as stated in Section 2, we construct a much larger dataset containing ten million training samples. In addition to the dataset from the previous section, we also added the following tasks:
After training the LLaMa-2-7b on this 10-million-sample dataset with the same experiment setup with Appendix A, we find the LLM exhibit a greater knowledge-forgetting but a promising performance in other tasks apart from knowledge benchmarks.
Task Name Baseline Result
NER 42.1 82.2
Program Execution 18.7 78.5
Toxic Classification 96 97.4
Question Generation 46.2 61.1
Text2sql 56 96.2
WSC 65.4 70.2
winogrande 61.7 66.1
Flores 0.1 26.0
Xsum 19.7 33.2
Race-middle 30.5 87.0
Race-high 30.4 83.3
RTE 52.7 87.4
ReCoRD 29.5 56.6
AX-g 52.0 87.9
multiRC 44.0 86.0
TriviaQA 52.2 30.9
NQ 18.5 14.2
Filtered TriviaQA 33.5 15.7
Filtered NQ 7.8 5.0
HotpotQA 11.2 7.6
Table 5: Performance of Llama-2-7B after vanilla SFT with a 10-million-sample datasets. There is a much more severe decrease in the performance on the CBQA tasks, while a great enhancement in other tasks compared with the baseline.
Appendix C Mixed Distribution Dilemmas for Expert Balancing
$L(\mathbf{X})=$ $\prod_{x\in\mathbf{X}_{1}}\left(m\mathcal{N}\left(x;\mu^{\prime}_{1},\sigma^{\prime 2}_{1}\right)\right.$
$\left.+(1-m)\mathcal{N}\left(x;\mu^{\prime}_{2},\sigma^{\prime 2}_{2}\right)\right)$
×$\times\prod_{x\in\mathbf{X}_{2}}\left(m\mathcal{N}\left(x;\mu^{\prime}_{1},\sigma^{\prime 2}_{1}\right)\right.$
$\left.+(1-m)\mathcal{N}\left(x;\mu^{\prime}_{2},\sigma^{\prime 2}_{2}\right)\right),$ (10)
Task Name
# Expert=8
# rank=4
# Expert=4
# rank=4
# Expert=6
# rank=8
# Expert=6
# rank=16
WSC 71.2 76.0 70.2 76.9
winogrande 69.8 56.0 69.5 70.9
Flores 25.0 25.8 26.1 26.3
Xsum 32.8 33.3 33.7 34.0
Race-middle 90.3 84.2 90.3 90.5
Race-high 87.1 80.7 87.3 87.2
RTE 84.5 80.1 88.1 85.2
ReCoRD 85.6 85.5 86.0 86.1
AX-g 88.8 77.5 88.2 85.7
multiRC 77.2 87.6 81.1 87.3
TriviaQA 54.4 57.8 58.2 58.9
NQ 25.6 27.9 27.8 28.2
Filtered TriviaQA 30.7 35.8 36.7 34.3
Filtered NQ 11.5 13.4 12.0 15.4
HotpotQA 14.5 16.0 16.4 16.5
Table 6: Detailed result on sensitivity study.
$\frac{\partial\log L(\mathbf{X})}{\partial\mu_{1}^{\prime}}=$ $\sum_{x\in\mathbf{X}_{1}\cup\mathbf{X}_{2}}\frac{\partial}{\partial\mu_{1}^{\prime}}\log\left(m\mathcal{N}_{1}(x)\right.$
$\left.+(1-m)\mathcal{N}_{2}(x)\right)$
$=$ $\sum_{x\in\mathbf{X}_{1}\cup\mathbf{X}_{2}}\left(\frac{x-\mu_{1}^{\prime}}{\sigma_{1}^{\prime 2}}\right)$
×$\times\frac{m\mathcal{N}_{1}(x)}{m\mathcal{N}_{1}(x)+(1-m)\mathcal{N}_{2}(x)},$ (11)
$\frac{d\log L}{dm}=$ $\frac{\partial\log L}{\partial\mu_{1}^{\prime}}\frac{d\mu_{1}^{\prime}}{dm}$
$=$ $\left(\sum_{x\in\mathbf{X}_{1}\cup\mathbf{X}_{2}}\left(\frac{x-\mu_{1}^{\prime}}{\sigma_{1}^{\prime 2}}\right)\right.$
×$\quad\left.\times\frac{m\mathcal{N}_{1}(x)}{m\mathcal{N}_{1}(x)+(1-m)\mathcal{N}_{2}(x)}\right)$
×$\quad\times\frac{p_{1}(\mu_{2}-\mu_{1})}{m^{2}}$
$\leq$ $0,$ (12)
The inverse result can be derived similarly. Therefore, the best training error is achieved only when the mixing coefficient $m$ of the prior distribution is consistent with the actual sampling distribution weight $p_{1}$.
Appendix D Detalied Results of Sensitivity Study
Table 6 shows the detailed results presented in Section 4.3.